A total of 1,025 tickets were sold for a game for a total of $1,250. If adult tickets sold for $2.00 and children's tickets sold for $1.00, how many of each kind of ticket were sold?
Answer by mananth(16946)

x= No of adult tickets
y= No of child tickets

1 x + 1 y = 1025 .............1

2 x + 1 y = 1250 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-1 x -1 y = -1025
2 x + 1 y = 1250
Add the two equations
1 x = 225
/ 1
x = 225
plug value of x in (1)
1 x + 1 y = 1025
225 + y = 1025
y = 1025 -225
y = 800
y = 800 1
x= 225 No of adult tickets
y= 800 No of child tickets
m.ananth@hotmail.ca